OpenStudy (anonymous):

5u+3v=-17 and 2u+v=-6

OpenStudy (anonymous):

\[5u+3v=-17\] \[2u+v=-6\] multiple 2nd eq by -3 and add thenm together

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

6u+3v=-6

OpenStudy (anonymous):

no, -6u+-3v=18 then add two eq together and what do you get?

OpenStudy (anonymous):

-u + v = 1

OpenStudy (anonymous):

no, -u=1 so u=-1, do you see your mistakes now take u=-1 and find v?

OpenStudy (anonymous):

The v was eliminated in the subtraction step.

OpenStudy (anonymous):

V=-5

OpenStudy (anonymous):

does v=-5 and u=-1 work for both equations

OpenStudy (anonymous):

you will find that v does not = -5 so check your math, actully so me your math and I'll help you understand your mistake.

OpenStudy (anonymous):

2(-1)+v=-7 -2+v=-7 -2+5=-7

OpenStudy (anonymous):

V=5

OpenStudy (anonymous):

look at your equation is not right because =-6 on -7 so try again

OpenStudy (anonymous):

2(-1)+v=-6 -2+v=-6 -2+-4=-6 V=-4

OpenStudy (anonymous):

good nice job

OpenStudy (anonymous):

thank you for your help